OpenStudy (anonymous):

How do I find the lowest common multiple for: 4m-8, 6m-18?

OpenStudy (anonymous):

4m-8=4(m-2), 6m-18=6(m-3) The lowest common multiple of 4 and 6 is 12. Thus the LCM of these two is 12*(m-2)*(m-3)